A rectangle is 5 times as long as it is wide. If the are is 125 square feet, find its perimeter?
Answer by checkley77(12844)

L=5W
AREA=LW
125=5W*W
125=5W^2
W^2=125/5
W^2=25
W=SQRT25
W=5 ANS. FOR THE WIDTH.
L=5*5=25 ANS. FOR THE LENGTH.